OpenStudy (anonymous):

Name two fractions that are equivalent to 3/4

myininaya (myininaya):

1=2/2 and 1=3/3 and so on... Correct? So what is (3/4)(2/2) and (3/4)(3/3)

OpenStudy (anonymous):

is it 6/8 and 9/12??

OpenStudy (anonymous):

yes 6/8 and 9/12 would be equivalent fractions to 3/4